Assuming a return on stocks of 13%, what is the value of a company that pays a current dividend of $3.20 and will grow at 6% for 2 years and then 4% forever? PLEASE TYPE YOUR ANSWER
Answer #1

D1=(3.2*1.06)=3.392

D2=(3.392*1.06)=3.59552

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(3.59552*1.04)/(0.13-0.04)

=41.5482311

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.392/1.13+3.59552/1.13^2+41.5482311/1.13^2

=$38.36(Approx).